OpenStudy (shamil98):

The accompanying table shows time-to-speed data for a 2004 Audi S4 Qauttro Cabriolet accelerating from rest to 130mph. How far had the Cabriolet traveled by the time it reached this speed? ( Use trapezoids to estimate the area under the velocity curve, but be careful: the time intervals vary in length.)

OpenStudy (shamil98):

--------------------------------------------------------- Speed Change: Time: Zero to (sec) 30 mph 2.0 40 mph 3.2 50 mph 4.5 60 mph 5.8 70 mph 7.7 80 mph 9.5 90 mph 11.6 100 mph 14.9 110 mph 17.8 120 mph 21.7 130 mph 26.3

OpenStudy (nincompoop):

have you tried diff methods of solving for the area yet?

OpenStudy (shamil98):

Nope, gotta stick with trapezoids for this.

OpenStudy (nincompoop):

so how does the trapezoidal rule work?

OpenStudy (shamil98):

\[T = \frac{ h }{ 2 } (y_0 + 2y_1 + 2y_2 +.... 2y_{n-1} + y_n)\]

OpenStudy (nincompoop):

then use it

OpenStudy (shamil98):

\[h = \frac{ b-a }{ n }\] n = the number of intervals.

OpenStudy (nincompoop):

i know what n stands for laughing out loud

OpenStudy (nincompoop):

\[\int\limits_{a}^{b}f(x) dx = (b-a)\left[ \frac{ (f(a)+f(b) }{ 2 } \right]\]

OpenStudy (shamil98):

\[T = \frac{ 50 }{ 11 } (2 + 2(3.2) + 2(4.5) + 2(5.8) + 2(7.7) + 2(9.5) + 2(11.6) +2(14.9) \] \[+ 2(17.8) + 2(21.7) + 26.3)\]

OpenStudy (shamil98):

\[T = \frac{ 50 }{ 11 } (250) = \frac{ 12500 }{11 }\] The answer from the book is 3340 ft. So this method is wrong ..

OpenStudy (nincompoop):

multiple trapezoidal rule

OpenStudy (timaashorty):

No... \[n= number of nuts you contain \]

OpenStudy (shamil98):

The equation that the book solution uses is totally different from what I just did.

OpenStudy (nincompoop):

applying trapezoidal rule on each segment follows \[\int\limits_{a}^{ b}f(x) dx =\frac{ b-a }{ 2n }\left[ f(a)+2\left\{ \sum_{i=1}^{n}f(a+ih) \right\}+f(b)\right]\]

OpenStudy (shamil98):

I haven't learned that formula yet -.-

OpenStudy (shamil98):

I'll prob just ask my teacher tomorrow thanks for helping nin.

OpenStudy (nincompoop):

or learn it on your own :P

OpenStudy (nincompoop):

if you can do the rectangles pretty well, I am sure trapezoid is not more difficult

OpenStudy (nincompoop):

what other rules have you learned if I may ask

OpenStudy (shamil98):

Simpsons/trapezoidal, rectangle, cylinder i think thats about it.